OpenStudy (anonymous):

Solve By using Substitution: 2x+y=-92 2x+2y=-98

OpenStudy (anonymous):

Please show work, i will give you a medal if you answer

OpenStudy (anonymous):

solve for y in the first equation: y = -92 - 2x then substitute this in for y in the second equation: 2x + 2(-92 - 2x) = -98 expand and simplify: 2x - 184 - 4x = -98 -2x = 86 x = -43 then substitute this in for x in the first equation: 2(-43) + y = -92 y = -92 + 86 = -6 your solution is (-43, -6)

OpenStudy (anonymous):

thankyou!